OpenStudy (anonymous):

16/120 simplest form

OpenStudy (inkyvoyd):

okay this one isn't hard: divide the top and bottom by 8

OpenStudy (anonymous):

16/8=2 120/2= 60

OpenStudy (inkyvoyd):

uhh 16/8=2 120/8=15

OpenStudy (inkyvoyd):

so your answer is \(\frac{2}{15}\)

OpenStudy (anonymous):

oh yeah sorry i devided it by 2 srry thx for catching me i wrote that down

OpenStudy (anonymous):

Do you know what the greatest common divisor is? In general you need to find the GCD of the numerator and denominator. In this case it is indeed 8.